# ForestSplats: Deformable transient field for Gaussian Splatting in the Wild

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has emerged, showing real-time rendering speeds and high-quality results in static scenes. Although 3D-GS shows effectiveness in static scenes, their performance significantly degrades in real-world environments due to transient objects, lighting variations, and diverse levels of occlusion. To tackle this, existing methods estimate occluders or transient elements by leveraging pre-trained models or integrating additional transient field pipelines. However, these methods still suffer from two defects: 1) Using semantic features from the Vision Foundation model (VFM) causes additional computational costs due to pre-training on external data to use prior knowledge. 2) The transient field requires significant memory to handle transient elements with per-view Gaussians and struggles to define clear boundaries for occluders, solely relying on photometric errors. To address these problems, we propose ForestSplats, a novel approach that leverages the deformable transient field and a superpixel-aware mask to efficiently represent transient elements in the 2D scene across unconstrained image collections and effectively decompose static scenes from transient distractors without VFM. We designed the transient field to be deformable, capturing per-view transient elements. Furthermore, we introduce a superpixel-aware mask that clearly defines the boundaries of occluders by considering photometric errors and superpixels. Additionally, we propose uncertainty-aware densification to avoid generating Gaussians within the boundaries of occluders during densification. Through extensive experiments across several benchmark datasets, we demonstrate that ForestSplats outperforms existing methods without VFM and shows significant memory efficiency in representing transient elements.
@@ 1 Introduction
Novel view synthesis from a sparse set of captured 2D images is a challenging and fundamental problem in computer vision and graphics, which has recently gained popularity. Such advancements play a critical role in applications such as augmented/virtual reality (VR) [6, 5], autonomous driving [12, 47], and 3D content generation [15, 36]. Even though Neural Radiance Field (NeRF) [18] has shown impressive results and rendering efficiency, it suffers from slow rendering speeds. Due to this issue, 3D Gaussian Splatting (3D-GS) [7] recently has emerged as an explicit 3D representation that offers real-time rendering speeds. However, existing methods [35, 31] assume that 2D images are captured without transient elements, various levels of occluders, or appearance variations such as changing sky, weather, and illumination. Therefore, accurately and robustly reconstructing with unconstrained image collections remains a primary challenge due to many transient elements in real-world environments. Several works have proposed various solutions to tackle this challenge. Early NeRF-based methods [17, 2, 38] introduce per-image transient embedding or photometric errors to typically ignore transient elements from unconstrained images. NeRF-W [17] optimizes an uncertainty embedding to address distractors. CR-NeRF [40] proposes a cross-ray paradigm and grid sampling strategy for efficient appearance modeling. Although showing promising results, these methods [17, 2, 40] are computationally intensive and struggle with real-time rendering. In contrast, 3D-GS-based methods [37, 10, 24, 14, 33] usually utilize semantic features from Vision Foundation Model (VFM) to estimate precise distractors mask, showing fast rendering speed. Wild-GS [37] uses depth information from Depth Anything [39] and a hierarchical decomposition strategy to regularize transient elements and preserve geometry consistency. WildGaussians [10] also utilizes semantic features from DINOv2 [20] to handle occlusions. HybridGS [14] defines a transient field to explicitly decompose transient elements. Despite significant advancements, these methods still suffer from two limitations: Firstly, leveraging semantic features from the Vision Foundation Model (VFM) causes additional computational costs due to pre-computation on external huge data to use prior knowledge. Secondly, the transient field uses significant memory to handle transient elements using Gaussians per image and struggles to clearly estimate boundaries of occluders due to solely relying on the photometric errors. Our motivation stems from the lack of appearance variation and multi-view consistency in transient elements. That is, we only need to consider transient elements in the 2D scene from each view of unconstrained image scenarios. To end, we propose ForestSplats, which leverages a deformable transient field and superpixel-aware mask to effectively decouple transient elements from static scenes, showing high-quality rendering. We design deformable transient field that efficiently represents transient elements in the 2D scene across unconstrained image collections, which are explicitly decomposed from static field and show memory efficiency. Furthermore, we introduce superpixel-aware masking, which effectively separates distractors from static scenes without a pre-trained VFM by considering the photometric errors and superpixels in the 2D scene. Additionally, we incorporate a multi-stage training scheme to enhance the superpixel-aware mask in accurately capturing transient elements, achieving distractor-free novel-view synthesis. Furthermore, we introduce uncertainty-aware densification for static field, thereby avoiding generating Gaussians within the boundaries of occluders, showing high-quality rendering. Through extensive experiments on several datasets, we demonstrate that ForestSplats effectively decomposes transient elements from static scenes, outperforming the state-of-the-art methods without VFM in novel view synthesis. Our primary contributions can be summarized as follows.
@@ 2 Related Works
Recent novel view synthesis methods [18, 7, 26] rely on the assumption of distractor-free environments. However, unconstrained image collections often violate these assumptions due to dynamic elements, lighting variations, and other transient factors. Numerous earlier methods [17, 18, 40, 11] utilize implicit neural representation to encode 3D scenes, with differentiable volume rendering along the rays. NeRF-W [17] optimizes a transient field and an uncertainty embedding per view to address distractors. CR-NeRF [40] introduces a cross-ray paradigm and a grid sampling strategy for efficient appearance modeling. NeRF-MS [11] tackles multi-sequence images captured in the wild using triplet loss and a decomposition module with several transient embeddings. While NeRF-based methods [17, 18, 40, 11] offer efficient and high-quality 3D rendering, they are still computationally resource-intensive and limit real-time rendering speed. On the other hand, 3D-GS-based methods [7, 39, 37, 10] explicitly represent the 3D scene with differentiable rasterization, showing fast rendering speed and high-quality results. Wild-GS [37] leverages depth information from pre-trained Depth Anything [39] and a hierarchical decomposition strategy to regularize transient elements and preserve geometry consistency. WildGaussians [10] also utilizes semantic features of pre-trained DINOv2 [20] to handle transient elements. Unlike these approaches, our method effectively decomposes distractors from static scenes and represents transient elements without pre-trained VFM by leveraging photometric errors and superpixels.
@@
Adaptive Density Control of 3D Gaussians. Adaptive density control (ADC) strategy in 3D Gaussian Splatting works [7] through two operations, densification and pruning, which capture empty areas or fine details during optimization. While showing efficiency, this ADC module relies on predetermined thresholds and does not leverage context information, leading to blurriness in high-frequency details in sparse point areas. To tackle this, Pixel-GS [46] leverages the number of pixels covered by Gaussian primitives as weights to promote densification, complementing average 2D position gradients. Taming-3DGS [16] proposes score-based densification and steerable densification strategies, ordered by high opacity, improving the quality per Gaussian while avoiding an overemphasis on specific regions during training. Unfortunately, most existing methods [7, 46, 16] improve rendering quality in constrained image collections, but their application to unconstrained image collections is limited and remains challenging due to transient elements. In this work, we aim to develop an efficient densification strategy for casually captured images in the wild. To this end, we assign learnable uncertainty attributes to each Gaussian, removing high-uncertainty Gaussians in the static field. Furthermore, we effectively refine densification to avoid generating Gaussians within the boundaries of occluders during densification by considering positional gradient and the number of pixels covered by Gaussians.
@@ 3 Preliminary
[…]
@@
Figure 3: Comparison of transient masking methods. (a) CR-NeRF focuses only on the photometric errors, (b) WildGaussians considers the semantic features, and (c) ForestSplats leverages photometric errors and superpixels to capture transient elements.
@@ 4.1 Deformable Transient Field
Given transient embedding per image $t_{i}$ and position $\mu_{i}$ of transient 3D Gaussians $\mathcal{G}_{d}(\mu_{i},s_{i},r_{i})$ as inputs, the deformation transient MLP $f_{d}$ predicts transient elements, producing the offset $\Delta$$\mu$, $\Delta$$s$, and $\Delta$$r$. Furthermore, transient elements in unconstrained images have inconsistent colors across views. Therefore, we additionally introduce the transient color MLP $f_{c}$ to map the color of transient elements expressed as:
@@ 4.2 Superpixel-aware Mask
$+\lambda_{0}\mathcal{L}_{BCE}(\mathcal{M}_{O},\mathcal{M}_{S})+\lambda_{1}||\mathcal{M}_{O}||^{2},$ (7)
where $\lambda_{0}$ is $23e-$ 2, and $\lambda_{1}$ is $45e-$ 4. We observe that initially incorporating the transient mask during training leads to an ambiguous separation between the static and transient fields. To this end, we propose a multi-stage training scheme to resolve the ambiguity between the static and transient fields. Initially, we train the static field to represent the entire image as follows:
@@
where ¯$\bar{\mathcal{M}}_{S}$ is calculated as ¯$\bar{\mathcal{M}}_{S}=1-\mathcal{M}_{S}$. Finally, we jointly train static and transient fields as shown in Eq. 4.2. This multi-stage training scheme enhances the static field to capture static scenes while achieving high-quality results.
@@
GS-W [44] 0.55 / 71 18.67 0.595 0.288 21.50 0.783 0.152 18.52 0.644 0.335 19.56 0.674 0.258
NeRF On-the-go [22] 43 / $<$1 20.63 0.661 0.191 22.31 0.780 0.130 22.19 0.753 0.169 21.71 0.731 0.163
WildGaussians [10] 0.50 / 108 20.62 0.658 0.235 22.80 0.811 0.092 23.03 0.771 0.172 22.15 0.756 0.167
SpotLessSplats [24] 0.34 / 93 20.01 0.596 0.276 22.79 0.816 0.161 21.91 0.71 0.222 21.57 0.707 0.219
DeSplat [33] 0.24 / 106 19.93 0.695 0.170 23.47 0.845 0.100 24.33 0.870 0.105 22.58 0.803 0.125
HybridGS [14] - 21.42 0.684 0.268 23.51 0.830 0.160 23.05 0.768 0.204 22.66 0.761 0.211
@@ 4.3 Uncertainty-aware Densification
where ^$\hat{\mathcal{M}}_{l^{(u)}}$ is calculated as $(\sum_{k=1}^{K}l_{k}^{(u)})\odot\mathcal{M}_{S}$. According to existing methods [46, 4], large Gaussians often cover a small area depending on the viewpoint. Inspired by this, we improve the splitting process by shifting static Gaussians along their positional gradient $\mathbf{g}_{i}$, weighted by the number of pixels covered by each Gaussian, computed as $\mathbf{g}_{i}\times(\frac{|\{p\mid p\in\mathcal{G}_{i}\}|}{H\times W})$. Subsequently, we apply random sampling from the PDF to avoid splitting the static Gaussians lying in a region of transient elements as shown in Fig. 6-(b). For above the threshold, similar to 3D-GS-based methods, we remove the static Gaussians with high uncertainty to ensure high-quality and consistent rendering results.
@@ 5 Experiments
11footnotetext: The results are reproduced from the official code.
@@
Following existing methods [10], we evaluate our method on two datasets: the NeRF On-the-go dataset [22] and the Photo Tourism dataset [27]. The NeRF On-the-go dataset [22] contains twelve scenes, casually captured in indoor and outdoor settings, with three varying levels of occlusion (ranging from 5% to 30%). For a fair comparison, we utilize six scenes and follow the version of the dataset where all images were undistorted as [10]. The Photo Tourism dataset [27] includes several scenes of well-known monuments. Each scene consists of a collection of user-uploaded unconstrained images captured at different dates, times of day, and exposure levels, covering a wide range of lighting conditions. In our experiments, we utilize three scenes: Brandenburg Gate, Sacre Coeur, and Trevi Fountain. We evaluate our method without including occluders in the image, following previous methods [40, 10, 37]. To demonstrate the effectiveness of our method, we use PSNR, SSIM [34], and LPIPS [45] as metrics to evaluate the performance of our proposed method.
@@
Table 3: Ablation of computational efficiency. We report CPU Memory usage (MB) for initial the number of transient Gaussians.
@@
Implementation details. We develop our ForestSplats method based on WildGaussians [10]. We follow the hyperparameter settings for the default setting and train using the Adam [8] optimizer without weight decay. Specifically, we conducted the improved ADC solely for a static field. Furthermore, we extended the implementation by adapting the Pixel-GS [46] to calculate the number of per-pixel Gaussians using a custom CUDA kernel. For transient elements, we adopt the transient field initialized by randomly selecting 10K points from those obtained through SfM. The deformable MLP consists of eight fully connected layers. In the initial training phase, we conduct 150K steps of training solely on the static field to represent the entire image. Subsequently, we proceed to joint training, where both static and transient fields are optimized until 200K iterations. Finally, we focus on training the static field to capture the static scene. For a fair comparison, we use NerfBaselines [9] as our evaluation framework, providing a unified interface to the originally released source codes. All experiments are performed on an NVIDIA RTX 24GB GPU.

### v3

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has emerged, showing real-time rendering speeds and high-quality results in static scenes. Although several 3D-GS methods show effectiveness in static scenes, their performance significantly degrades in real-world environments due to transient objects, lighting variations, and diverse levels of occlusion. To tackle this, existing methods estimate occluders or transient elements by leveraging pre-trained models or integrating additional transient field pipelines. However, these methods still suffer from two defects: 1) Using semantic features from the Vision Foundation Model (VFM), such as DINO, might limit generalization on unseen data due to reliance on prior knowledge. 2) The transient field requires significant memory to handle transient elements with per-view Gaussians and struggles to define clear boundaries for occluders, solely relying on photometric errors. To address these problems, we propose ForestSplats, a novel approach that leverages the deformable transient field and a superpixel-aware mask to efficiently represent transient elements in the 2D scene across unconstrained image collections and effectively decompose static scenes from transient distractors without VFM. We designed the transient field to be deformable, capturing per-view transient elements. Furthermore, we introduce a superpixel-aware mask that clearly defines the boundaries of occluders by considering photometric errors and superpixels. Additionally, we propose uncertainty-aware densification to avoid generating Gaussians within the boundaries of occluders during densification. Through extensive experiments across several benchmark datasets, we demonstrate that ForestSplats outperforms existing methods without VFM and shows significant memory efficiency in representing transient elements.
@@
footnotetext: [Uncaptioned image] Page: pull-ups/ForestSplats    [Uncaptioned image] Model: pull-ups/Collections
@@ 1 Introduction
Novel view synthesis from a sparse set of captured 2D images is a challenging and fundamental problem in computer vision and graphics, which has recently gained popularity. Such advancements play a critical role in applications such as augmented/virtual reality (VR) [5, 4], autonomous driving [11, 45], and 3D content generation [13, 34]. Even though Neural Radiance Field (NeRF) [16] has shown impressive results and rendering efficiency, it suffers from slow rendering speeds. Due to this issue, 3D Gaussian Splatting (3D-GS) [6] recently has emerged as an explicit 3D representation that offers real-time rendering speeds. However, existing methods [33, 29] assume that 2D images are captured without transient elements, various levels of occluders, or appearance variations such as changing sky, weather, and illumination. Therefore, accurately and robustly reconstructing with unconstrained image collections remains a primary challenge due to many transient elements in real-world environments.
Several works have proposed various solutions to tackle this challenge. Early NeRF-based methods [15, 1, 36] introduce per-image transient embedding or photometric errors to typically ignore transient elements from unconstrained images. CR-NeRF [38] proposes a cross-ray paradigm and grid sampling strategy for efficient appearance modeling. Although showing promising results, these methods [15, 1, 38] are computationally intensive and struggle with real-time rendering. In contrast, 3D-GS-based methods [35, 9, 22, 12, 31] usually utilize semantic features from Vision Foundation Model (VFM) to estimate precise distractors mask, showing fast rendering speed. WildGaussians [9] also utilizes semantic features from DINOv2 [18] to handle occlusions. HybridGS [12] defines a transient field to explicitly decompose transient elements. Despite significant advancements, these methods still suffer from two limitations: Firstly, leveraging semantic features from the Vision Foundation Model (VFM) rely on prior knowledge, which restricts effective generalization to unseen data. Secondly, the transient field uses significant memory to handle transient elements using Gaussians per image and struggles to clearly estimate boundaries of occluders due to solely relying on the photometric errors.
Our motivation stems from the lack of appearance variation and multi-view consistency in transient elements. That is, we only need to consider transient elements in the 2D scene from each view of unconstrained image scenarios. To end, we propose ForestSplats, which leverages a deformable transient field and superpixel-aware mask to effectively decouple transient elements from static scenes, showing high-quality rendering. We design deformable transient field that efficiently represents transient elements in the 2D scene across unconstrained image collections, which are explicitly decomposed from static field and show memory efficiency. Furthermore, we introduce superpixel-aware masking, which effectively separates distractors from static scenes without a pre-trained VFM by considering the photometric errors and superpixels in the 2D scene. Additionally, we incorporate a multi-stage training scheme to enhance the superpixel-aware mask in accurately capturing transient elements, achieving distractor-free novel-view synthesis. Furthermore, we introduce uncertainty-aware densification for static field, thereby avoiding generating Gaussians within the boundaries of occluders, showing high-quality rendering. Through extensive experiments on several datasets, we demonstrate that ForestSplats effectively decomposes transient elements from static scenes, outperforming the state-of-the-art methods without VFM in novel view synthesis. Our primary contributions can be summarized as follows.
@@ 2 Related Works
Uncertainty in 3D Scene Reconstruction. Recent novel view synthesis methods [16, 6, 24] rely on the assumption of distractor-free environments. However, unconstrained image collections often violate these assumptions due to dynamic elements, lighting variations, and other transient factors. Numerous earlier methods [15, 16, 38, 10] utilize implicit neural representation to encode 3D scenes, with differentiable volume rendering along the rays. NeRF-W [15] optimizes a transient field and an uncertainty embedding per view to address distractors. CR-NeRF [38] introduces a cross-ray paradigm and a grid sampling strategy for efficient appearance modeling. NeRF-MS [10] tackles multi-sequence images captured in the wild using triplet loss and a decomposition module with several transient embeddings. While NeRF-based methods [15, 16, 38, 10] offer efficient and high-quality 3D rendering, they are still computationally resource-intensive and limit real-time rendering speed. On the other hand, 3D-GS-based methods [6, 37, 35, 9] explicitly represent the 3D scene with differentiable rasterization, showing fast rendering speed and high-quality results. Wild-GS [35] leverages depth information from pre-trained Depth Anything [37] and a hierarchical decomposition strategy to regularize transient elements and preserve geometry consistency. WildGaussians [9] also utilizes semantic features of pre-trained DINOv2 [18] to handle transient elements. Unlike these approaches, our method effectively decomposes distractors from static scenes and represents transient elements without pre-trained VFM by leveraging photometric errors and superpixels.
@@
Adaptive Density Control of 3D Gaussians. Adaptive density control (ADC) strategy in 3D Gaussian Splatting works [6] through two operations, densification and pruning, which capture empty areas or fine details during optimization. While showing efficiency, this ADC module relies on predetermined thresholds and does not leverage context information, leading to blurriness in high-frequency details in sparse point areas. To tackle this, Pixel-GS [44] leverages the number of pixels covered by Gaussian primitives as weights to promote densification, complementing average 2D position gradients. Taming-3DGS [14] proposes score-based densification and steerable densification strategies, ordered by high opacity, improving the quality per Gaussian. Unfortunately, most existing methods [6, 44, 14] improve rendering quality in constrained image collections, but their application to unconstrained image collections remains challenging due to transient elements. In this work, we aim to develop an efficient densification strategy for casually captured images in the wild. To this end, we assign learnable uncertainty attributes to each Gaussian, removing high-uncertainty Gaussians in the static field. Furthermore, we effectively refine densification to avoid generating Gaussians within the boundaries of occluders during densification by considering positional gradient and the number of pixels covered by Gaussians.
@@ 3 Preliminary
3D Gaussian Splatting (3D-GS) [6] has emerged as a promising method for novel view synthesis, representing a scene as a set of differentiable Gaussians $\{\mathcal{G}_{i}\}$. Each Gaussian $\mathcal{G}_{i}$ is parameterized by a position $\mu_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, which is decomposed into rotation matrix $R\in\text{SO(3)}$, and scaling matrix $S\in\mathbb{R}^{3\times 3}$, an opacity $\alpha_{i}\in[0,1]$, and view-dependent colors $c_{i}\in\mathcal{C}^{N_{sh}}$ represented via spherical harmonics (SH) coefficients $N_{sh}$. To render images, 3D Gaussians $\{\mathcal{G}_{i}\}$ are projected onto screen space as 2D Gaussians via the viewing transformation $W\in\mathbb{R}^{3\times 3}$ and covariance matrix $\Sigma^{\prime}_{i}=JW\Sigma_{i}W^{T}J^{T}$ where $J\in\mathbb{R}^{2\times 3}$ is the Jacobian of the Taylor approximation of the projective transformation. For each pixel, the $K$ Gaussians $\{\mathcal{G}_{k}\;|\;k=1,...,K\}$ are ordered by depth and rendered using alpha-blending, resulting in pixel color $\hat{C}$:
@@
Figure 3: Comparison of masking methods. (a) CR-NeRF focuses only on the photometric errors such as L1 loss, (b) WildGaussians considers the semantic features, relying on VFM, and (c) ForestSplats leverages photometric errors and superpixels to capture distractors, creating sharp boundaries of transients without VFM.
@@ 4.1 Deformable Transient Field
[…]
@@ 4.2 Superpixel-aware Mask
[…]
where $\lambda_{0}$ is 3e23e-2, and $\lambda_{1}$ is 5e45e-4. We observe that initially incorporating the transient mask during training leads to an ambiguous separation between the static and transient fields. To this end, we propose a multi-stage training scheme to resolve the ambiguity between the static and transient fields. Initially, we train the static field to represent the entire image as follows:
@@
where $\bar{\mathcal{M}}_{S}$ is calculated as $\bar{\mathcal{M}}_{S}=1-\mathcal{M}_{S}$. Finally, we jointly train static and transient fields as shown in Eq. 4.2. This multi-stage training scheme enhances the static field to capture static scenes while achieving high-quality results.
@@
GS-W [42] 1.03 / 105 18.67 0.595 0.288 21.50 0.783 0.152 18.52 0.644 0.335 19.56 0.674 0.258
NeRF On-the-go [20] 43 / <<1 20.63 0.661 0.191 22.31 0.780 0.130 22.19 0.753 0.169 21.71 0.731 0.163
WildGaussians [9] 1.56 / 94 20.62 0.658 0.235 22.80 0.811 0.092 23.03 0.771 0.172 22.15 0.756 0.167
SpotLessSplats [22] 1.87 / 81 20.01 0.596 0.276 22.79 0.816 0.161 21.91 0.710 0.222 21.57 0.707 0.219
DeSplat [31] 0.84 / 106 19.93 0.695 0.170 23.47 0.845 0.100 24.33 0.870 0.105 22.58 0.803 0.125
HybridGS [12] 0.98 / 112 21.42 0.684 0.268 23.51 0.830 0.160 23.05 0.768 0.204 22.66 0.761 0.211
@@ 4.3 Uncertainty-aware Densification
where $\hat{\mathcal{M}}_{l^{(u)}}$ is calculated as $(\sum_{k=1}^{K}l_{k}^{(u)})\odot\mathcal{M}_{S}$. According to existing methods [44, 3], large Gaussians often cover a small area depending on the viewpoint. Inspired by this, we improve the splitting process by shifting static Gaussians along their positional gradient $\mathbf{g}_{i}$, weighted by the number of pixels covered by each Gaussian, computed as $\mathbf{g}_{i}\times(\frac{|\{p\mid p\in\mathcal{G}_{i}\}|}{H\times W})$. Subsequently, we apply random sampling from the PDF to avoid splitting the static Gaussians lying in a region of transient elements as shown in Fig. 6-(b). For above the threshold, similar to 3D-GS-based methods, we remove the static Gaussians with high uncertainty to ensure high-quality and consistent rendering results.
@@ 5 Experiments
Following existing methods [9], we evaluate our method on two datasets: the NeRF On-the-go dataset [20] and the Photo Tourism dataset [25]. The NeRF On-the-go dataset [20] contains twelve scenes, casually captured in indoor and outdoor settings, with three varying levels of occlusion (ranging from 5% to 30%). For a fair comparison, we utilize six scenes and follow the version of the dataset where all images were undistorted as [9]. The Photo Tourism dataset [25] includes several scenes of well-known monuments. Each scene consists of a collection of user-uploaded unconstrained images captured at different dates, times of day, and exposure levels, covering a wide range of lighting conditions. In our experiments, we utilize three scenes: Brandenburg Gate, Sacre Coeur, and Trevi Fountain. We evaluate our method without including occluders in the image, following previous methods [38, 9, 35]. To evaluate our method, we use PSNR, SSIM [32], and LPIPS [43] as metrics.
@@
Table 3: Ablation of computational efficiency. We report memory usage (MB) for initial the number of transient Gaussians.
@@
Implementation details. We develop our ForestSplats method based on WildGaussians [9]. We follow the hyperparameter settings for the default setting and train using the Adam [7] optimizer without weight decay. Specifically, we conducted the improved ADC solely for a static field. Furthermore, we extended the implementation by adapting the Pixel-GS [44] to calculate the number of per-pixel Gaussians using a custom CUDA kernel. For transient elements, we adopt the transient field initialized by randomly selecting 10K points from those obtained through SfM. In the initial training phase, we conduct 150K steps of training solely on the static field. Subsequently, we proceed to joint training, where both static and transient fields are optimized until 200K iterations. Finally, we focus on training the static field to capture the static scene. All experiments are performed on an NVIDIA RTX 24GB GPU using NerfBaselines [8] as our evaluation framework for a fair comparison.